A 44mH inductor is connected to 220v, 50Hz ac supply. Determine the rms value of the current in the circuit
1
Expert's answer
2020-04-17T10:15:56-0400

The inductive reactance

"X_L=2\pi f L=2\pi \times 50\times 44\times 10^{-3}=13.8\:\rm \Omega"

The rms value of the current


"I_{rms}=\frac{V_{rms}}{X_L}=\frac{220}{13.8}=15.9\:\rm A."